A toddler starts running down a sidewalk at 230 feet per minute. One minute later, a worried mother runs after the child at 660 feet per minute. When will the mother overtake the toddler?
Answer by rwm(914)

230(x+1)=660x
230x+230=660x
230=430x
divide by 10
23/43=x